# Frustratingly Easy Test-Time Adaptation of Vision-Language Models

### v1

@@ Abstract
Vision-Language Models seamlessly discriminate among arbitrary semantic categories, yet they still suffer from poor generalization when presented with challenging examples. For this reason, Episodic Test-Time Adaptation (TTA) strategies have recently emerged as powerful techniques to adapt VLMs in the presence of a single unlabeled image. The recent literature on TTA is dominated by the paradigm of prompt tuning by Marginal Entropy Minimization, which, relying on online backpropagation, inevitably slows down inference while increasing memory. In this work, we theoretically investigate the properties of this approach and unveil that a surprisingly strong TTA method lies dormant and hidden within it. We term this approach Zero (TTA with “zero” temperature), whose design is both incredibly effective and frustratingly simple: augment $N$ times, predict, retain the most confident predictions, and marginalize after setting the Softmax temperature to zero. Remarkably, Zero requires a single batched forward pass through the vision encoder only and no backward passes. We thoroughly evaluate our approach following the experimental protocol established in the literature and show that Zero largely surpasses or compares favorably w.r.t. the state-of-the-art while being almost 10×10$\times$10 × faster and 13×13$\times$13 × more memory friendly than standard Test-Time Prompt Tuning. Thanks to its simplicity and comparatively negligible computation, Zero can serve as a strong baseline for future work in this field.
@@ 2.2 How does MEM affect the marginal probability distribution?
The recent literature on TTA shows that minimizing $\mathcal{L}_{ent}$ significantly enhances the robustness of model outputs. However, the impact of this process on the marginal probability distribution $\overline{p}$ remains unclear. We start with a straightforward hypothesis: due to its nature, minimizing $\mathcal{L}_{ent}$ tends to increase the probability of the already prevailing class of $\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})$. More formally, denoting with $\hat{c}$ the prediction of $\overline{p}$ (i.e., $\hat{c}=\operatorname*{arg\,max}\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})$), we hypothesize that $\overline{p}(y=\hat{c}|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}}^{*})>\overline{p}(y=\hat{c}|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})$. If this hypothesis is realized, it comes as a natural consequence that minimizing $\mathcal{L}_{ent}$ is unlikely to alter the prevailing class of $\overline{p}$, thus resulting in a consistent prediction pre- and post-TTA where $\operatorname*{arg\,max}\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})=\operatorname*{arg\,max}\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}}^{*})$.
@@ 2.3 How does $\overline{p}$ relate to the standard inference protocol?
From prior work on Test-Time Augmentations (TTAug) with unimodal neural networks [39, 34], empirical evidence suggests that $\overline{p}(\cdot|\mathbf{x})$ is more robust than $p(\cdot|\mathbf{x})$. This observation leads to the hypothesis that the expected risk of predicting with $\overline{p}$ is lower than that of doing so with $p$. However, the literature lacks guarantees for this hypothesis, except for the peculiar case in which the risk function is the squared error, i.e., $\ell(a,b)=(a-b)^{2}$ [15]. [^1] In Appendix C, we show that this bound generalizes to any function $\ell$ satisfying the triangular inequality.
@@
Revisiting model ensembling for TTA. Eq. (6) holds as long as all events modeled as Bernoulli processes are independent. Thus, we have an equivalent error estimate for the setup in which only a single classifier $f$ is present and $X_{y}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ is a set of independent examples with the same underlying label $y$. Within this framework, any group of $k$ examples in $X_{y}$ to which the classifier has assigned the same label $\hat{y}$ is also a set of independent Bernoulli processes, whose error probability is still quantified via Eq. (6). Note that this resembles the TTA setup in the presence of $N$ views of the source sample $\mathbf{x}$, as they all share the same label and are only dependent through $\mathbf{x}$.
$\overline{p}$ is better than $p$ (if $f$ is calibrated). The final step can be taken through the lens of model calibration [7], a property requiring that the confidence of a classifier matches its accuracy. For example, a classifier $f$ whose confidence is $0.7$ is expected to be correct $70\%$ of the times, and viceversa, if $f$ is calibrated. In the previous discussion, if we denote with $k(y)$ the number of examples correctly labeled as $y$, then the accuracy of the classifier is exactly $k(y)/N$. It follows that $\overline{p}(y)\approx k(y)/N$ if $f$ exhibits good calibration and that the probability of picking the wrong class with this marginal probability is still quantified by Eq. (6). Given this relationship, we have that $\overline{p}(y)=\max\overline{p}(\cdot)$ if $k(y)$ matches or exceeds the majority within $N$. Thus, the probability of picking the wrong class with $\overline{p}$ is approximated by marginalizing out all values of $k$ that satisfy this criterion, which entails that the error or $\overline{p}$ can be expressed with the cumulative distribution of (6):
@@ 3.1 Augmentations undermine the reliability of $\overline{p}$
Augmentations lead to poor calibration. We demonstrate the impact of augmentation-induced overconfidence using the same model and datasets of Section 2.3. For each dataset, we generate an augmented counterpart following the augmentation and filtering setup of TPT [37], i.e.: we augment an input $N=64$ times using simple random resized crops and horizontal flips. Then, we only retain 10% of the $N$ views according to confidence-based filtering, resulting in 6 views per sample. Consequently, each augmented dataset contains 6×6$\times$6 × more data points than its plain counterpart. The Expected Calibration Error (ECE) [7] reported in Appendix B conveys that ① zero-shot CLIP is well-calibrated (ECE${}<0.1$ for all datasets), strongly supporting the theory of Section 2.3 and ② the augmented visual space greatly increases the calibration error.
Poor calibration is always caused by overconfidence. We investigate the reason for the increase in ECE by presenting reliability diagrams for the ImageNet validation set in Fig. 1(b). In a reliability diagram, every bar below the identity line $y=x$ signals overconfidence (i.e., the confidence on the x-axis prevails over the accuracy on the y-axis), while the opposite signals under-confidence. Notably, the ECE increases exclusively due to overconfidence. The error rate, in contrast, decreases slightly. As displayed in Appendix B, this phenomenon persists across all datasets.
@@ 4 Experiments
Method ImageNet A V2 R Sketch Average
@@
In this section, we present a comprehensive experimental evaluation of Zero. Our results show that Zero, alongside its simplicity, is an effective and efficient approach for TTA.
@@ 4.1 Experimental Protocol
Group 2: When comparing to PromptAlign, we follow Samadh et al. [32] and start from a MaPLe initialization for a fair comparison. Within this group, we also report TPT on top of MaPLe, as in [32].
@@
Group 3: When comparing to RLCF, we use both CLIP-ViT-B-16 and CLIP-ViT-L-14 as in [53]. Specifically, confidence-based filtering acts on top of the output of the first model, and the selected inputs are passed to the second model for the final output. Both forward passes are inevitable in RLCF, so this scheme corresponds to “early-exiting” the pipeline, exactly as per MEM. RLCF can vary according to (i) the parameter group being optimized and (ii) the number of adaptation steps. We denote with $\Theta_{v}$ the full image encoder tuning, with $\bm{\mathbf{t}_{\text{ctx}}}$ prompt tuning, and with $t$ the number of adaptation steps. For example, RLCF ${}_{t=3}^{\bm{\mathbf{t}_{\text{ctx}}}}$ indicates RLCF with prompt tuning for 3 TTA steps.
@@
Benchmarks. We follow the established experimental setup of [37, 32], evaluating Zero on Natural Distribution Shifts and Fine-grained Classification. For the former, we consider the ImageNet validation set and the four datasets for Natural Distribution Shifts already presented in Section 2, commonly considered Out-of-Distribution (OOD) datasets for CLIP. For fine-grained classification, we evaluate all TTA methods on 9 datasets. Specifically, we experiment with Oxford-Flowers (FWLR) [24], Describable Textures (DTD) [2], Oxford-Pets (PETS) [28], Stanford Cars (CARS) [16], UCF101 (UCF) [40], Caltech101 (CAL)[5], Food101 (FOOD) [1], SUN397 (SUN)[47], and FGVC-Aircraft (AIR) [22] [^2] We find that the extremely OOD domain of satellite imagery (i.e., EuroSAT [10]) leads to consistent failures of all TTA methods. Hence, we dedicate an entire section to the study of this dataset in Appendix E.. For these, we refer to the test split in Zhou et al. [55] as per the common evaluation protocol. Unless otherwise specified, all tables in this section report top-1 accuracy.
@@
Implementation Details. $\mathcal{A}$ only contains random resized crops and random horizontal flips. We do not tune any hyperparameter; instead, we inherit the setup of TPT with $N=64$ and the top percentile for confidence-based filtering set to $0.1$. To ensure hardware differences do not play any role in comparisons, we execute all TTA methods under the same hardware setup by running the source code of each repository with no modifications. We always use 1 NVIDIA A100 GPU and FP16 Automatic Mixed Precision. Results are averaged over 3 different seeds.
Table 2: Finegrained classification. TTA methods are grouped according to the reference baseline, top-1 accuracy is reported and bold text indicates the best performer of each group.
Method FLWR DTD PETS CARS UCF CAL FOOD SUN AIR Avg
CLIP-ViT-B-16
Zero-Shot 67.44 44.27 88.25 65.48 65.13 93.35 83.65 62.59 23.67 65.98
Ensemble 67.07 45.09 88.28 66.16 67.51 93.91 84.04 66.26 23.22 66.84
TPT 68.75 47.04 87.23 66.68 68.16 93.93 84.67 65.39 23.13 67.22
Zero 67.07 45.80 86.74 67.54 67.64 93.51 84.36 64.49 24.40 66.84
Zero+Ensemble 66.82 45.86 87.20 68.48 68.57 94.14 84.58 66.90 24.42 67.44
MaPLe
Zero-Shot 72.23 46.49 90.49 65.57 68.69 93.53 86.20 67.01 24.74 68.33
TPT 72.37 45.87 90.72 66.50 69.19 93.59 86.64 67.54 24.70 68.57
PromptAlign 72.39 47.24 90.76 68.50 69.47 94.01 86.65 67.54 24.80 69.04
Zero 71.20 47.70 90.17 67.91 69.49 94.12 86.78 67.55 25.57 68.94
CLIP-ViT-B-16 + CLIP-ViT-L-14
ZeroShot 75.76 51.83 92.86 76.16 73.70 94.04 88.03 66.96 30.54 72.21
RLCF ${}_{t=1}^{\bm{\mathbf{t}_{\text{ctx}}}}$ 71.58 50.34 89.01 69.76 69.84 94.09 85.90 67.33 23.71 69.06
RLCF ${}_{t=3}^{\bm{\mathbf{t}_{\text{ctx}}}}$ 72.49 51.93 89.55 72.91 72.31 95.00 86.84 69.04 25.40 70.61
RLCF ${}_{t=1}^{\Theta_{v}}$ 72.56 52.21 89.51 63.12 71.49 94.65 86.90 68.50 24.06 69.22
RLCF ${}_{t=3}^{\Theta_{v}}$ 71.74 53.27 91.15 70.93 73.24 94.73 87.28 69.38 28.54 71.14
Zero 75.34 54.22 92.90 77.33 74.26 94.52 87.57 68.05 32.11 72.92
@@ 4.2 Results
Group 1 (TPT): Zero surpasses TPT consistently on all datasets. Among OOD datasets, peak difference is reached with ImageNet-A, where Zero outperforms TPT by $+6.58\%$. Enriching Zero with hand-crafted prompts improves results further, with an average margin of $+3.80\%$ w.r.t. TPT.
@@
Group 2 (PromptAlign): Within the second comparison group, Zero outperforms PromptAlign on all datasets, with $+1.89\%$ being the gap in average performance. Note also that Zero consistently outperforms TPT also when the baseline initialization is MaPLe (by an average of $+3.13\%$).
@@
Group 3 (RLCF): We follow [53] and report RLCF variants with $t=3$ steps. In this group, Zero outperforms RLCF in 3 out of 5 datasets, with a gap in the average performance of $+0.99\%$. Importantly, RLCF is only close to Zero with image encoder tuning; only prompt tuning is insufficient.
@@
Fine-grained Classification. Results for fine-grained classification are shown in Table 2. To foster readability, the standard deviations of Zero are separately reported in Table 6 (Appendix).
@@
Group 1 (TPT): Default Zero improves over the zero-shot baseline CLIP-ViT-B-16, but is slightly outperformed by TPT with an average margin of $-0.38\%$. However, extending Zero with hand-crafted prompts (something that TPT cannot do by design) is sufficient to outperform TPT on 5 out of 9 datasets, and obtain an average improvement of $+0.22\%$.
@@
Group 2 (PromptAlign): On average, PromptAlign has a slight improvement of $+0.1\%$ over Zero. However, note that this is mostly influenced by the performance on one dataset only (Oxford-Flowers) and that, in contrast, Zero surpasses PromptAlign in 6 out of 9 datasets. In line with the previous benchmark, Zero better adapts MaPLe than TPT, outperforming it in 7 out of 9 datasets.
@@
Group 3 (RLCF): As Zhao et al. [53] do not report results on fine-grained classification, we use their code to evaluate four RLCF variants: $\Theta_{v}$ and $\bm{\mathbf{t}_{\text{ctx}}}$ tuning, with $t=1$ and $t=3$ adaptation steps. We find that Zero largely outperforms RLCF regardless of the configuration. Even with respect to the strongest RLCF ${}_{t=3}^{\Theta_{v}}$ variant, Zero obtains an average improvement of $+1.78\%$.
@@
Computational Requirements. The complexity of Zero does not scale linearly with the size of the label space, as it does for prompt-tuning strategies. To quantify the computational gain of Zero w.r.t. other TTA methods, we report the runtime per image and peak GPU memory in Table 3 under the same hardware requirements (1 NVIDIA RTX 4090). We compare to TPT and the RLCF pipeline in a worst-case scenario where the label space is large (ImageNet). We omit PromptAlign from our analysis since it has slightly worse computational performance than TPT.

### v2

@@ Abstract
Vision-Language Models seamlessly discriminate among arbitrary semantic categories, yet they still suffer from poor generalization when presented with challenging examples. For this reason, Episodic Test-Time Adaptation (TTA) strategies have recently emerged as powerful techniques to adapt VLMs in the presence of a single unlabeled image. The recent literature on TTA is dominated by the paradigm of prompt tuning by Marginal Entropy Minimization, which, relying on online backpropagation, inevitably slows down inference while increasing memory. In this work, we theoretically investigate the properties of this approach and unveil that a surprisingly strong TTA method lies dormant and hidden within it. We term this approach Zero (TTA with “zero” temperature), whose design is both incredibly effective and frustratingly simple: augment $N$ times, predict, retain the most confident predictions, and marginalize after setting the Softmax temperature to zero. Remarkably, Zero requires a single batched forward pass through the vision encoder only and no backward passes. We thoroughly evaluate our approach following the experimental protocol established in the literature and show that Zero largely surpasses or compares favorably w.r.t. the state-of-the-art while being almost 10×10$\times$10 × faster and 13×13$\times$13 × more memory friendly than standard Test-Time Prompt Tuning. Thanks to its simplicity and comparatively negligible computation, Zero can serve as a strong baseline for future work in this field. The code is available.
@@ 2.2 How does MEM affect the marginal probability distribution?
The recent literature on TTA shows that minimizing $\mathcal{L}_{ent}$ significantly enhances the robustness of model outputs. However, the impact of this process on the marginal probability distribution $\overline{p}$ remains unclear. We start with a straightforward hypothesis: due to its nature, minimizing $\mathcal{L}_{ent}$ tends to increase the probability of the most probable class of $\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})$. More formally, denoting with $\hat{c}$ the prediction of $\overline{p}$ (i.e., $\hat{c}=\operatorname*{arg\,max}\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})$), we hypothesize that $\overline{p}(y=\hat{c}|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}}^{*})>\overline{p}(y=\hat{c}|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})$. If this hypothesis is realized, it comes as a natural consequence that minimizing $\mathcal{L}_{ent}$ is unlikely to alter the prevailing class of $\overline{p}$, thus resulting in a consistent prediction pre- and post-TTA where $\operatorname*{arg\,max}\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}})=\operatorname*{arg\,max}\overline{p}(\cdot|\mathbf{x},\bm{\mathbf{t}_{\text{ctx}}}^{*})$.
@@ 2.3 How does $\overline{p}$ relate to the standard inference protocol?
From prior work on Test-Time Augmentations (TTAug) with unimodal neural networks [40, 35], empirical evidence suggests that $\overline{p}(\cdot|\mathbf{x})$ is more robust than $p(\cdot|\mathbf{x})$. This observation leads to the hypothesis that the expected risk of predicting with $\overline{p}$ is lower than that of doing so with $p$. However, the literature lacks guarantees for this hypothesis, except for the peculiar case in which the risk function is the squared error, i.e., $\ell(a,b)=(a-b)^{2}$ [16]. [^1] In Appendix D, we show that this bound generalizes to any function $\ell$ satisfying the triangular inequality.
@@
Revisiting model ensembling for TTA. Eq. (6) holds as long as all events modeled as Bernoulli processes are independent. Thus, we have an equivalent error estimate for the setup in which only a single classifier $f$ is present and $X_{y}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ is a set of independent examples with the same underlying label $y$. Within this framework, any group of $k$ examples in $X_{y}$ to which the classifier has assigned the same label $\hat{y}$ is also a set of independent Bernoulli processes, whose error probability is still quantified via Eq. (6). Note that this resembles the TTA setup in the presence of $N$ views of the source sample $\mathbf{x}$, as long as augmentations do not change their underlying labels. We refer the reader to Appendix H for a discussion about the independence assumption among different views.
$\overline{p}$ is better than $p$ (if $f$ is calibrated). The final step can be taken through the lens of model calibration [8], a property requiring that the confidence of a classifier matches its accuracy. For example, a calibrated classifier $f$ whose confidence is $0.7$ is expected to be correct $70\%$ of the times. In the previous discussion, if we denote with $k(y)$ the number of examples correctly labeled as $y$, then the accuracy of the classifier is exactly $k(y)/N$. It follows that there is a positive correlation between accuracy and confidence if $f$ exhibits good calibration, i.e., $\uparrow k(y)/N\implies\uparrow\overline{p}(y)$. Thus, the probability of picking the wrong class with this marginal probability is approximated by Eq. (6). Given this relationship, we have that $\overline{p}(y)=\max\overline{p}(\cdot)$ if $k(y)$ matches or exceeds the majority within $N$. Thus, the probability of picking the wrong class with $\overline{p}$ is approximated by marginalizing out all values of $k$ that satisfy this criterion, which entails that the error or $\overline{p}$ can be expressed with the cumulative distribution of (6):
@@ 3.1 Augmentations undermine the reliability of $\overline{p}$
Augmentations lead to poor calibration. We demonstrate the impact of augmentation-induced overconfidence using the same model and datasets of Section 2.3. For each dataset, we generate an augmented counterpart following the augmentation and filtering setup of TPT [38], i.e.: we augment an input $N=64$ times using simple random resized crops and horizontal flips. Then, we only retain 10% of the $N$ views according to confidence-based filtering, resulting in 6 views per sample. Consequently, each augmented dataset contains 6×6$\times$6 × more data points than its plain counterpart. The Expected Calibration Error (ECE) [8] reported in Appendix C conveys that ① zero-shot CLIP is well-calibrated (ECE${}<0.1$ for all datasets), strongly supporting the theory of Section 2.3 and ② the augmented visual space greatly increases the calibration error.
Poor calibration is frequently linked to overconfidence. We investigate the reason for the increase in ECE by presenting reliability diagrams for the ImageNet validation set in Fig. 1(b). In a reliability diagram, every bar below the identity line $y=x$ signals overconfidence (i.e., the confidence on the x-axis prevails over the accuracy on the y-axis), while the opposite signals under-confidence. Notably, in the scope of our experiments, overconfidence is the primal factor leading to an increase in the ECE. The error rate, in contrast, decreases slightly. In Appendix C, we also experiment across all datasets for Natural Distribution Shifts and different CLIP models pretrained on the 2B subset of LAION [2, 34]. Importantly, this phenomenon further persists within this extended experimental suite.
@@ 4 Experiments
Method ImageNet A V2 R Sketch Mean
@@
Zero 75.52$\pm$0.03 75.15$\pm$0.26 70.37$\pm$0.05 87.21$\pm$0.09 59.61$\pm$0.04 73.57
In this section, we present a comprehensive experimental evaluation of Zero. Similarly to [38, 33, 54], we always work in the setup of single test point adaptation. Our results show that Zero, alongside its simplicity, is an effective and efficient approach for TTA.
@@ 4.1 Experimental Protocol
Group 2: When comparing to PromptAlign, we follow Samadh et al. [33] and start from a MaPLe initialization for a fair comparison. MaPLe prompts are learned on ImageNet, following [33]. Within this group, we also report TPT on top of MaPLe, as in [33].
@@
Group 3: When comparing to RLCF, we use both CLIP-ViT-B-16 and CLIP-ViT-L-14 as in [54]. Specifically, confidence-based filtering acts on top of the output of the first model, and the selected inputs are passed to the second model for the final output. Both forward passes are inevitable in RLCF, so this scheme corresponds to “early-exiting” the pipeline, exactly as per MEM. RLCF can vary according to (i) the parameter group being optimized and (ii) the number of adaptation steps. We denote with $\Theta_{v}$ the full image encoder tuning, with $\bm{\mathbf{t}_{\text{ctx}}}$ prompt tuning, and with $t$ the number of adaptation steps. For example, RLCF ${}_{t=3}^{\bm{\mathbf{t}_{\text{ctx}}}}$ indicates RLCF with prompt tuning for 3 TTA steps. Note that, since all methods need to forward more than one image to the teacher model, the zero-shot baseline of this group is exactly zero-shot classification with CLIP-ViT-L-14.
Pretrainings. This Section deals with models officially released by OpenAI [28]. Appendix B further reports experiments with LAION-pretrained CLIP models [2], as well as the soft prompt initialization with supervised Context Optimization (CoOp) from [56].
@@
Benchmarks. We follow the established experimental setup of [38, 33], evaluating Zero on Natural Distribution Shifts and Fine-grained Classification (also referred to as “Cross-Datasets Generalization" in previous works). For the former, we consider the ImageNet validation set and the four datasets for Natural Distribution Shifts already presented in Section 2, commonly considered Out-of-Distribution (OOD) datasets for CLIP. For fine-grained classification, we evaluate all TTA methods on 10 datasets. Specifically, we experiment with Oxford-Flowers (FWLR) [25], Describable Textures (DTD) [3], Oxford-Pets (PETS) [29], Stanford Cars (CARS) [17], UCF101 (UCF) [41], Caltech101 (CAL)[6], Food101 (FOOD) [1], SUN397 (SUN)[48], FGVC-Aircraft (AIR) [23] and EuroSAT (ESAT) [11]. For all of these datasets, we refer to the test split in Zhou et al. [56] as per the common protocol.
Textual prompts. When +Ensemble is specified, we do not use dataset-specific templates. In contrast, we use the set of 7 generic templates highlighted in the official CLIP repository [28] across all datasets. When adapting MaPLe, we stick to the ImageNet-learned prompts released by [15] and evaluate them cross-datasets as in [33].
@@
Implementation Details. The augmentation pool $\mathcal{A}$ only contains random resized crops and random horizontal flips. The only hyperparameter of Zero is the percentile for confidence-based filtering, which is set to $0.3$ after validation on ImageNet (following standard practice [51]) and kept fixed for all datasets. We inherit the setup of TPT with $N=64$, crafting $63$ augmentations to collate with the source image. To ensure hardware differences do not play any role in comparisons, we execute all TTA methods under the same hardware setup by running the source code of each repository with no modifications. We always use 1 NVIDIA A100 GPU and FP16 Automatic Mixed Precision. Results are averaged over 3 different seeds. Unless otherwise specified, all tables report top-1 accuracy.
@@ 4.2 Results
Group 1 (TPT): Zero surpasses TPT consistently on all datasets. Among OOD datasets, peak difference is reached with ImageNet-A, where Zero outperforms TPT by $+4.84\%$. Enriching Zero with hand-crafted prompts improves results further, with an average margin of $+3.66\%$ w.r.t. TPT.
@@
Group 2 (PromptAlign): Within the second comparison group, Zero outperforms PromptAlign on all datasets, with $+1.68\%$ being the gap in average performance. Zero consistently outperforms TPT also when the baseline initialization is MaPLe (by an average of $+2.92\%$). Please note that we omit evaluation on ImageNet for this group, since PromptAlign adopts token-level statistics from this dataset when adapting to test points, which would render the comparison unfair. For completeness, we report that zero-shot MaPLe achieves an accuracy of $70.72\%$ on ImageNet, which is improved to $72.99\%$ by adapting with Zero ($+2.27\%$).
@@
Group 3 (RLCF): We follow [54] and report RLCF variants with $t=3$ steps. In this group, Zero outperforms RLCF in 5 out of 5 datasets, with a gap in the average performance of $+1.25\%$. Importantly, RLCF is only close to Zero with image encoder tuning; only prompt tuning is insufficient.
Fine-grained Classification.
Table 2: Fine-grained classification. TTA methods are grouped according to the reference baseline, top-1 accuracy is reported and bold text indicates the best performer of each group.
Method FLWR DTD PETS CARS UCF CAL FOOD SUN AIR ESAT Mean Median
CLIP-ViT-B-16
Zero-Shot 67.44 44.27 88.25 65.48 65.13 93.35 83.65 62.59 23.67 42.01 63.58 65.31
Ensemble 67.07 45.09 88.28 66.16 67.51 93.91 84.04 66.26 23.22 50.42 65.20 66.66
TPT 68.75 47.04 87.23 66.68 68.16 93.93 84.67 65.39 23.13 42.86 64.78 67.42
Zero 67.68 46.12 87.75 68.04 67.77 93.66 86.53 65.03 25.21 34.33 64.21 67.72
Zero+Ensemble 67.17 45.86 87.83 68.97 69.18 94.41 86.77 67.63 25.21 42.17 65.52 68.30
MaPLe
Zero-Shot 72.23 46.49 90.49 65.57 68.69 93.53 86.20 67.01 24.74 48.06 66.30 67.85
TPT 72.37 45.87 90.72 66.50 69.19 93.59 86.64 67.54 24.70 47.80 66.49 68.36
PromptAlign 72.39 47.24 90.76 68.50 69.47 94.01 86.65 67.54 24.80 47.86 66.92 68.99
Zero 71.62 47.89 90.60 68.58 69.87 94.48 87.20 68.20 26.25 39.47 66.42 69.23
CLIP-ViT-B-16 + CLIP-ViT-L-14
Zero-Shot 75.76 51.83 92.86 76.16 73.70 94.04 88.03 66.96 30.54 54.38 70.43 74.73
RLCF ${}_{t=1}^{\bm{\mathbf{t}_{\text{ctx}}}}$ 71.58 50.34 89.01 69.76 69.84 94.09 85.90 67.33 23.71 46.87 66.84 69.80
RLCF ${}_{t=3}^{\bm{\mathbf{t}_{\text{ctx}}}}$ 72.49 51.93 89.55 72.91 72.31 95.00 86.84 69.04 25.40 45.96 68.14 72.40
RLCF ${}_{t=1}^{\Theta_{v}}$ 72.56 52.21 89.51 63.12 71.49 94.65 86.90 68.50 24.06 47.74 67.07 70.00
RLCF ${}_{t=3}^{\Theta_{v}}$ 71.74 53.27 91.15 70.93 73.24 94.73 87.28 69.38 28.54 47.41 68.77 71.34
Zero 76.41 53.63 94.08 78.39 74.68 95.21 90.66 69.61 33.62 44.21 71.05 75.55
@@
Results for fine-grained classification are shown in Table 2. To foster readability, the standard deviations of Zero are separately reported in Table 11 (Appendix).
@@
Group 1 (TPT): Default Zero improves over the zero-shot baseline CLIP-ViT-B-16, but is outperformed by TPT with an average margin of $-0.57\%$. However, extending Zero with hand-crafted prompts (something that TPT cannot do by design) is sufficient to outperform TPT on 7 out of 10 datasets, and obtain an average improvement of $+0.74\%$.
@@
Group 2 (PromptAlign): On average, PromptAlign has an improvement of $+0.5\%$ over Zero. However, note that this is mostly influenced by the performance on one dataset only (EuroSAT) and that, in contrast, Zero surpasses PromptAlign in 7 out of 10 datasets. In line with the previous benchmark, Zero better adapts MaPLe than TPT, again outperforming it in 7 out of 10 datasets.
@@
Group 3 (RLCF): As Zhao et al. [54] do not report results on fine-grained classification, we use their code to evaluate four RLCF variants: $\Theta_{v}$ and $\bm{\mathbf{t}_{\text{ctx}}}$ tuning, with $t=1$ and $t=3$ adaptation steps. We find that Zero largely outperforms RLCF regardless of the configuration. Even with respect to the strongest RLCF ${}_{t=3}^{\Theta_{v}}$ variant, Zero obtains an average improvement of $+2.28\%$.
@@
Computational Requirements. The complexity of Zero does not scale linearly with the size of the label space, as it does for prompt-tuning strategies. To quantify the computational gain of Zero w.r.t. other TTA methods, we report the runtime per image and peak GPU memory in Table 3 under the same hardware (i.e., 1 NVIDIA RTX 4090). We compare the computational requirements of Zero to TPT and the RLCF pipeline in a worst-case scenario where the label space is large (ImageNet). We omit PromptAlign from our analysis since it has slightly worse computational performance than TPT.